Assuming that stars radiate like black bodies, the total radiation (energy) from a star depends on its... A) Temperature and Size B) Density and Size C) Size and distance from Earth D) Temperature and Mass E) Mass and Size

please explain reasoning or assumptions for the answer
Answer #1

Total energy density of a black body at temperature T,is
  u\propto T^4\Rightarrow u=AT^4
where, A is a constant.
So, the total energy is
  U=uV=AVT^4
So, the total energy depends upon the volume of the object, i.e., the size of the object and upon its temperature T.
So, option (A) is correct.
i.e.,
Temperature and size.